OpenStudy (anonymous):

Consider the proportion 3:4 = 6:8. What is true about the cross products?

OpenStudy (anonymous):

that the product of extremes is equal to the product of the means...soo product of means is 4*6=24 and product of extremes =8*3=24...

OpenStudy (jman):

cross products are equal